OpenStudy (anonymous):

What is the slope of the equation 4x + 8y = 24 a 1/3 b -1/3 c 1/2 d -1/2

OpenStudy (anonymous):

you have to subtract 4x from both sides first

OpenStudy (anonymous):

8y = 20?

OpenStudy (anonymous):

8y= -4x +24 , you can't subtract a number with a variable from a regular number. Then you have to divide each number by 8.. See what you come up with

OpenStudy (anonymous):

y = -1/2x + 3???????

OpenStudy (anonymous):

then do you subtract 3?

OpenStudy (anonymous):

You leave the 3, that's the y-intercept. Slope form is y=Mx+B. M= slope B= y-int. SO, your slope is -1/2

OpenStudy (anonymous):

ok thank you!!

OpenStudy (anonymous):

Your welcome! (: